OpenStudy (anonymous):

Find tan (A - B) for sin A = 12/13 with 0 < A < π/2 and cos B = 8/17 with 0 < B < π/2

OpenStudy (anonymous):

I would appreciate any help on these review questions: http://imgur.com/eBAAHd2 We're studying for the upcoming final and I'm super nervous! Help would be much appreciated as I could really use some!

OpenStudy (anonymous):

I'm tying to figure this out by using the pythagorean theorem with the sum and difference identities, but I'm having a rough time figuring it out.

OpenStudy (usukidoll):

hmm.. I think you need one of the identities.

OpenStudy (usukidoll):

but I really need to study for my own final and eat dinner... sorry :(

OpenStudy (usukidoll):

it has something to do with the identities... that's all I can throw out there

OpenStudy (campbell_st):

well for the given information of sin(A) = 12/13 draw a triangle  find x using pythagoras' theorem.... then you can find tan(A) you need to provide some information on B hope it helps

OpenStudy (campbell_st):

and you need another triangle for B  find y using pythagoras' theorem... then you can find tan B and the expansion of \[\tan(A - B) = \frac{\tan(A) - \tan(B)}{1 + \tan(A)\tan(B)}\] so substiute your ratios and evaluate

OpenStudy (anonymous):

@campbell_st Thanks! I figured that would be the formula we would use. If it would help any, my teacher provided us with this video: http://my.hrw.com/math11/math06_07/nsmedia/lesson_videos/alg2/player.html?contentSrc=7833/7833.xml But our examples are with ratio's and degrees: http://imgur.com/eBAAHd2 However I think I may be able to turn them into degrees.